OpenStudy (deathslasher99):

Simplify and write answer in standard form: (7f)(pv + 17f)

OpenStudy (lord_box):

You just have to distribute and then order the terms in decreasing power.

OpenStudy (deathslasher99):

Can you just give me the answer please i could really use it

OpenStudy (faiqraees):

We are not allowed to give answers according to OpenStudy policy

OpenStudy (deathslasher99):

well can you help me then?

OpenStudy (lord_box):

First, distribute through the parentheses

OpenStudy (deathslasher99):

I dont know non of this please explain better

OpenStudy (deathslasher99):

You act like if i knew what you where talking about

OpenStudy (deathslasher99):

I am not good at math

OpenStudy (deathslasher99):

This is why i would like some help

OpenStudy (deathslasher99):

Please

OpenStudy (lord_box):

One of the properties of multiplication is distributive, for instance, x(y + z) = xy + xz. Do this with the problem

OpenStudy (phi):

multiply each "term" inside the parens by 7f (7f)(pv + 17f) to multiply, write one term then * then the other term can you try ?

OpenStudy (faiqraees):

=a(b+c) =ab+ac where a = 7f, b= pv, c=17f

OpenStudy (skullpatrol):

recall: $$\Huge a(b+c) = ab + ac$$

OpenStudy (skullpatrol):

$$\Huge a(b+c) = ab + ac$$ $$\Huge (7f)(pv + 17f)= (7f)(pv)+(7f)(17f)$$ $$\Huge a(b+c) = ab + ac$$

OpenStudy (skullpatrol):

Simplify and write answer in standard form: $$\Huge (7f)(pv)+(7f)(17f)$$

OpenStudy (skullpatrol):

7 • 17 = ? $$f • f = f^2$$

OpenStudy (studyloverhojea):

(7f)(pv + 17f) do the distributive property 7f(pv) + 7f(17f) = 7fpv + 119f The answer is 7fpv + 119f. I think

OpenStudy (skullpatrol):

Don't forget the last term should be 119f^2